micpic234:

If ∠J measures 45°, ∠M measures 40°, and j is 7 feet, then find m using the Law of Sines. Round your answer to the nearest tenth. 6.4 ft 7.7 ft 8.5 ft 23.7 ft

mw104969:

which one do yuu think it is love??

micpic234:

8.5?

mw104969:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @micpic234 8.5? \(\color{#0cbb34}{\text{End of Quote}}\) no and yuu can also cancel out 6.4

micpic234:

ok is it 7.7 then becouse the 23.7 does not make sense

mw104969:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @micpic234 ok is it 7.7 then becouse the 23.7 does not make sense \(\color{#0cbb34}{\text{End of Quote}}\) actually it is 23.7

micpic234:

oh ok thanks

mw104969:

np love

jigglypuff314:

NO WAIT

jigglypuff314:

23.7 is incorrect

mw104969:

how come??

jigglypuff314:



jigglypuff314:

Here is the original set up ^ then the law of sines is: j / sin(J) = m / sin(M)

jigglypuff314:

You can set up the equation: 7 / sin(45) = m / sin(40) and solve for m

jigglypuff314:

@micpic234 come back, 23.7 was incorrect

kinglegend:

no thats wrong use a scientifc cacultor and plug in the formula

kinglegend:

😂 ez

jhonyy9:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jigglypuff314 You can set up the equation: 7 / sin(45) = m / sin(40) and solve for m \(\color{#0cbb34}{\text{End of Quote}}\) multiplie both sides by sin40 and look so what will get ? @Laylalyssa